OpenStudy (anonymous):

an automobile with an initial speed of 4.3 m/s accelerates uniformly at the rate of 3 m/s^2. find the final speed and the displacement after 5 s?

OpenStudy (abhisar):

use the equation v=u+at where, v=final speed u=initial speed a=acceleration t=time taken Find final velocity by using this equation and then use s=ut+\(\sf \frac{1}{2}at^2\) to find displacement

OpenStudy (perl):

@Abhisar can you help me

OpenStudy (perl):

v= 4.3 + 3*5

OpenStudy (perl):

final speed = 19.3 m/s displacement = 4.3*5 + 1/2( 3) (5^2) = 59